Derive an expression for the maximum potential energy of the pendulum bob of mass m in terms of the length, L and angular displacement, θ of the swinging pendulum.
1
Expert's answer
2020-09-17T14:29:41-0400

From the conservation of emergy:


"K_{max}=W=mgh=mg(L-L\cos{\theta})\\=mgL(1-\cos{\theta})"